a gas has a density of 1.98 g/L at standard temperature and pressure what is the molar mass
1
Expert's answer
2020-04-27T01:46:38-0400

At STP molar volume is

"V_m = 22.4 dm^3\/mol"

"density = \frac{M}{V_m}"

"M = density\times V_m = 1.98 \frac{g}{L}\times 22.4 \frac{L}{mol}= 44.4 \frac{g}{mol}"